### v2

@@ 2 Data
We present descriptive statistics showing the differences in the same-sex couple population compared to the different-sex couple population broken down by education level and sex in Table 3. Descriptively, there is a clear difference in labor market outcomes and characteristics between men and women in same-sex partnerships and those in different-sex partnerships when controlling for education.
@@ 3 Identification Strategy & Estimation
The differential roll-out of anti-discrimination laws by state and locality over time and between same-sex and different-sex couples lends itself to a saturated triple difference-in-differences framework with the main outcomes of interest as the change in labor supply and pay. This strategy uses the variation presented in Figure 1, which shows how the laws changed over time by state and county. Formally, we estimate this equation on labor supply:
@@
$Y_{ipct}$ $=\alpha\;+\;\beta_{1}\bigl{[}\mathrm{SSP}_{ipct}\times\mathrm{ENDA}_{ct}\bigr{]}\;+\;\sum_{s=1}^{S}\lambda_{s}\bigl{[}\mathrm{SSP}_{ipct}\times\mathrm{SSM}_{st}\times\mathbf{1}_{\text{state}(c)=s}\bigr{]}$ (1)
$\quad+\theta_{ct}+\mu_{pc}+\kappa_{pt}+X_{ipct}^{\prime}\,\delta+\varepsilon_{ipct}.$
@@
In our regression, equation (1) defines the labor market outcomes $Y_{ipct}$ for individual $i$, in partnership type $p$, residing in county $c$, at time $t$. These outcomes include labor force participation, employment status, and the percentile rank of real wages (in 2019 dollars). The binary variable $\mathrm{SSP}_{ipct}$ indicates whether an individual is in a same‐sex partnership (1) or a different‐sex partnership (0). $\mathrm{SSM}_{st}$ indicates whether state $s$ recognizes same‐sex marriage in year $t$, and $\mathbf{1}_{\{\mathrm{state}(c)=s\}}=1$ if county $c$ lies in state $s$. The error term $\epsilon_{ipct}$ is clustered at the county level to account for within-county correlation. [^12] The ACS only identifies metro counties, so the county fixed effect is close to a pseudo-county fixed effect. Rural counties within a state are compared together with a true county fixed effect for metro area counties. Our key variable of interest is the interaction of same‐sex partnership status with the presence of a county‐level Employment Non‐Discrimination Act, $\mathrm{ENDA}_{ct}$. [^13] Treatment status is computed using the earliest year of ENDA implementation at the county or state level, with individuals remaining classified as “treated” thereafter.
@@
We include three sets of high‐dimensional fixed effects to account for a wide range of unobserved factors. First, $\theta_{ct}$ (county×$\times$×year) absorbs any time‐varying shocks or policy changes common to all individuals within a county in a given year. Second, $\mu_{pc}$ (partnership×$\times$×county) controls for any unobserved, time‐invariant characteristics that are specific to each partnership type within a county. Third, $\kappa_{pt}$ (partnership×$\times$×year) accounts for year‐specific shocks that affect each partnership type differentially. The vector $X_{ipct}^{\prime}\delta$ includes additional demographic controls (e.g., race, age, and education). The coefficient of interest, $\beta_{1}$, identifies how ENDA coverage $\mathrm{ENDA}_{ct}$ differentially affects same‐sex partners, net of any confounding from same‐sex marriage legalization. The summation term $\sum_{s=1}^{S}\lambda_{s}\bigl{[}\mathrm{SSP}_{ipct}\times\mathrm{SSM}_{st}\bigr{]}$ partials out state‐specific labor‐market impacts of marriage legalization on same‐sex couples.
The key variation used to estimate $\beta_{1}$ is the differential timing of when counties adopted ENDAs, and how this interacted with the labor market outcomes for individuals in same-sex partnerships relative to those in different-sex partnerships. This variation allows us to isolate the effect of anti-discrimination laws on labor market outcomes by comparing same-sex and different-sex couples across counties with and without ENDAs, before and after their passage. Thus, $\beta_{1}$ captures the differential impact of ENDA on labor supply and wage outcomes specifically for individuals in same-sex partnerships. [^14] We recognize the heterogeneous impact (e.g., Burn (2018) of anti-discrimination laws across different regions, yet our analysis aims to estimate an average effect for those directly subject to the laws.
We limit our sample to people aged 25 to 65, and we estimate this model separately by sex using the ACS person weights. Following Chen and Roth (2024), we compute the percentile rank of annual and hourly real wages. [^15] Given the ACS’s 23-month reference period for earnings, as discussed by Hansen et al. (2020), this extended timeframe might not precisely align with the implementation dates of ENDAs, potentially biasing the immediate effects of these policies towards zero.
To interpret $\beta_{1}$ as the causal effect of enacting employment protections for sexual orientation, we assume that no other factors are driving changes in the same-sex versus different-sex wage differential in states that do and do not have LGBT employment protections. One concern is that the distribution of workers in same-sex partnerships is not random, and states/counties that pass anti-discrimination laws have a higher concentration of workers in same-sex partnerships. However, with county-year fixed effects, comparisons are made among individuals within the same county in the same year, ensuring that workers in San Francisco are compared to others in San Francisco in that same year. A potential concern would be if there is sorting of workers in same-sex partnerships to areas with anti-discrimination laws protecting them, changing the composition of the area. To account for this, we remove individuals who moved in the past year as a robustness check. [^16] While we address potential sorting by excluding recent movers, unobserved migration patterns could still affect the composition of same-sex partnerships in jurisdictions with anti-discrimination laws.
Another potential concern is that same-sex marriage legalization may confound our estimates. We address this by explicitly partialing out any state-specific effects of marriage legalization for same-sex partners. In particular, we include the term $\sum_{s=1}^{S}\lambda_{s}\bigl{[}\mathrm{SSP}_{ipct}\times\mathrm{SSM}_{st}\times\mathbf{1}_{\text{state}(c)=s}\bigr{]}$, which absorbs any differential labor-market impacts associated with the varying timing of same-sex marriage legalizations across states. This ensures that our key coefficient, $\beta_{1}$, isolates the independent effect of the local Employment Non-Discrimination Act. While these controls mitigate potential bias, we caution that, despite our efforts, overlapping policies may still exert residual confounding influences on the estimated effects.
To address potential biases from staggered policy adoption, we employ the methodology of Borusyak et al. (2021). Unlike standard TWFE models, which may misattribute post-treatment effects from early-treated units to later-treated ones, their approach uses only untreated observations to infer counterfactual outcomes for treated units. This strategy minimizes contamination and yields a more reliable estimate of $\beta_{1}$. Our identification hinges on the parallel trends assumption, that in the absence of anti-discrimination laws, labor market trends in counties with and without such laws would have evolved similarly. [^17] The difference-in-differences strategy hinges on the parallel trends assumption across groups and periods, with no anticipation of the laws prior to their implementation. We assess this assumption using event studies that estimate the dynamic effects of the policy, allowing us to observe pre-treatment trends and validate the plausibility of our identification strategy.
We report the coefficients estimated via Borusyak et al. (2021) in the main text using a fully saturated triple-difference (DDD) specification, and standard TWFE along with other difference-in-differences (DiD) estimators in Section 5. As we show in Section 5, the estimates for same-sex and different-sex individuals estimated separately are similar across estimators.

### v3

@@ 2 Data
We present descriptive statistics showing the differences in the same-sex couple population compared to the different-sex couple population broken down by education level and sex in Table 3. Descriptively, there is a clear difference in labor market outcomes and characteristics between men and women in same-sex couples and those in different-sex couples when controlling for education.
Table 1: Timing of Sexual Orientation Anti-discrimination Laws
Year State City or County
2005 and Before CA, CT, Boulder, CO; Denver, CO; Fort Collins, CO; Gainsville, FL; Hialeah, FL; Hollywood, FL;
DC, HI, Key West, FL; Miami, FL; Orlando, FL; Pembroke, FL; Saint Petersburg, FL;
ME, MD, Tampa, FL; West Palm, FL; Atlanta, GA; Ames, IA; Cedar Rapids, IA; Davenport, IA;
MA, MN, Des Moines, IA; Iowa City, IA; Campaigne, IL; Chicago, IL; Peoria, IL; Urbana, IL;
NV, NH, Bloomington, IN; Fort Wayne, IN; Michigan City, IN; Terre Haute, IN; Lawerence, KS;
NJ, NM, Covington, KY; Lexington, KY; Louisville, KY; New Orleans, LA; Ann Arbor, MI;
NY, RI, Detroit, MI; Grand Rapids, MI; Ypsilanti, MI; Columbia, MO; Kansas City, MO;
VT, WI Saint Louis, MO; Cleveland, OH; Columbus, OH; Toledo, OH; Eugene, OR;
Portland, OR; Benton County, OR; Salem, OR; Allentown, PA; Erie, PA; Harrisburg, PA;
Lancaster, PA; Philadelphia, PA; Pittsburgh; Scranton, PA; Austin, TX; Dallas, TX;
Fort Worth, TX; Alexandria, VA; Arlington, VA; Seattle, WA; Spokane, WA; Tacoma, WA
2006 IL, WA Dubuque, IA; Indianapolis, IN, Ferndale, MI; Lansing, MI;
Cincinnati, OH; Easton, PA; West Chester, PA; Charleston, SC
2007 CO, IA, OR Waterloo, IA; Coshocton, OH; Dayton, OH; Newark, OH; Charleston, WV
2008 Columbia, SC
2009 DE Allegheny, PA; Reading, PA; Salt Lake City, UT
2010 Tallahassee, FL; Traverse City, MI; Missoula, MT;
Lower Merion, PA; Grand County, UT; Summit County, UT
2011 Volusia County, FL; Evansville, IN; University City, MO; East Cleveland, OH,
Bethlehem, PA; Conshohocken, PA; Haverford, PA; Ogden, UT
2012 St. Augustine, FL; Boise, ID; New Albany, IN; South Bend, IN; Flint, MI;
Muskegon, MI; Maplewood, MO; Helena, MT; Omaha, NE; Canton, OH;
Abington, PA; Cheltenham, PA; Morgantown, WV
2013 Phoenix, AZ; Pocatello, ID; Frankfort, KY; Shreveport, LA; Battle Creek, MI;
Bristol, PA; Pittston, PA; San Antonio, TX; Charlottesville, VA; Huntington, WV
2014 Tempe, AZ; Adrian, MI; Macomb County, MI; Butte, MT; Atlantic Beach, FL
2015 UT Anchorage, AK; Osceola County, FL; Anderson, IN; Clinton, IN; Hammond, IN; Muncie, IN
2016 Kokomo, IN; Manahattan, KS; St. Charles, MO; Jackson, MS; Lakewood, OH;
Carlisle, PA; Dickson City, PA; Wilkes-Barre, PA; Martinsburg, WV; Wheeling, WV
2017 Sitka, AK; Glendale, AZ;
2018
2019 Fort Lauderdale, FL; Fernandina Beach, FL; Decatur, GA
2020 VA
  • Notes: List of states, cities, and counties with sexual orientation anti-discrimination laws pulled from LGBTMap.org, an advocacy group, as well as through media reports and local FOIA requests. We only list city or county laws if there is no state law.
Table 2: Counties with Largest LGB Populations
County State Percent of SSPs
San Francisco County CA 7.37
District of Columbia DC 7.20
New York County NY 5.43
Suffolk County MA 4.37
St. Louis city MO 4.01
Multnomah County OR 3.99
Alexandria city VA 3.70
Santa Fe County NM 3.43
Baltimore city MD 3.37
  • Notes: Using ACS and person weights to recover the percentage of partnerships that are same-sex partnerships by county over 2005-2019.
Table 3: Summary Statistics
Panel A: Men
High School Grad or Lower Some College or Higher
SSP DSP SSP DSP
n = 21,737 n = 2,805,327 n = 80,456 n = 4,943,301
Variable Mean Mean Difference Mean Mean Difference
In Labor Force 0.738 0.829 -0.091∗∗∗ 0.863 0.898 -0.035∗∗∗
(0.440) (0.377) (0.344) (0.303)
Employed 0.687 0.783 -0.096∗∗∗ 0.831 0.871 -0.041∗∗∗
(0.464) (0.412) (0.375) (0.335)
Real Wage 28,991.2 36,602.8 -7,611.6∗∗∗ 71,867.9 79,338.7 -7470.8∗∗∗
(40,363.5) (39,963.0) (89,316.6) (90,257.4)
Real Hourly Wage 16.735 18.661 -1.925∗∗∗ 35.718 37.691 -1.974∗∗∗
(58.837) (44.692) (76.987) (115.923)
Age 46.074 47.440 -1.367∗∗∗ 45.609 46.786 -1.177∗∗∗
(10.932) (10.912) (10.645) (11.071)
Number of Children 0.484 1.187 -0.702∗∗∗ 0.234 1.124 -0.890∗∗∗
(1.014) (1.280) (0.692) (1.187)
Asian 0.029 0.029 -0.000 0.046 0.062 -0.017∗∗∗
(0.169) (0.169) (0.209) (0.242)
Black 0.079 0.080 -0.001 0.041 0.057 -0.016∗∗∗
(0.269) (0.271) (0.198) (0.232)
Hispanic 0.115 0.118 -0.003 0.069 0.050 0.020∗∗∗
(0.319) (0.323) (0.254) (0.218)
White 0.789 0.798 -0.009∗∗ 0.863 0.842 0.021∗∗∗
(0.408) (0.401) (0.344) (0.365)
Panel B: Women
High School Grad or Lower Some College or Higher
SSP DSP SSP DSP
n = 22,238 n = 2,654,209 n = 79,414 n = 5,464,738
Variable Mean Mean Difference Mean Mean Difference
In Labor Force 0.724 0.596 0.128∗∗∗ 0.862 0.749 0.114∗∗∗
(0.447) (0.491) (0.345) (0.434)
Employed 0.673 0.558 0.115∗∗∗ 0.833 0.722 0.111∗∗∗
(0.469) (0.497) (0.373) (0.448)
Real Wage 25,032.0 16,555.5 8,476.5∗∗∗ 56,374.5 39,254.0 17,120.6∗∗∗
(33,315.4) (24,328.6) (64,852.8) (50,889.8)
Real Hourly Wage 14.331 10.497 3.834∗∗∗ 29.083 22.721 6.363∗∗∗
(37.390) (39.943) (62.691) (86.710)
Age 45.437 48.151 -2.714∗∗∗ 44.860 45.382 -0.522∗∗∗
(11.356) (10.973) (11.001) (11.134)
Number of Children 0.716 1.088 -0.371∗∗∗ 0.497 1.103 -0.606∗∗∗
(1.093) (1.287) (0.902) (1.179)
Asian 0.024 0.045 -0.021∗∗∗ 0.029 0.067 -0.038∗∗∗
(0.152) (0.206) (0.168) (0.250)
Black 0.107 0.065 0.042∗∗∗ 0.057 0.056 0.001
(0.309) (0.247) (0.231) (0.230)
Hispanic 0.099 0.120 -0.021∗∗∗ 0.058 0.054 0.004∗∗∗
(0.299) (0.325) (0.234) (0.226)
White 0.778 0.799 -0.022∗∗∗ 0.861 0.836 0.025∗∗∗
(0.416) (0.400) (0.346) (0.370)
  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019. Summary statistics are presented by education level. T-tests were conducted to determine significant differences between those in same-sex partnerships and different-sex partnerships. $p<0.10$, ∗∗ $p<0.05$, $^{{}^{***}}$ $p<0.01$.
((a)) 2005
((b)) 2019
Figure 1: Sexual Orientation and Anti-Discrimination Laws
Notes: State and local sexual orientation employment anti-discrimination laws in 2005 (a) and 2019 (b). Data on laws obtained from LGBTMap.org and authors’ own investigation using media reports and FOIA requests.
Figure 2: Percentage of Same Sex Couples: 2005–2019
Notes: Percentage of couples that are same-sex for each state and DC over 2005–2019 using the American Community Survey. Author’s calculations. District of Columbia’s share (7.14%) is a significant outlier relative to the other states.
@@ 3 Identification Strategy & Estimation
The differential roll-out of anti-discrimination laws by state and locality over time, and between same-sex and different-sex couples, naturally lends itself to a saturated triple-differences framework, where the primary outcomes of interest are changes in labor supply and pay. This strategy uses the variation presented in Figure 2, which shows how the laws changed over time by state and county. Formally, we estimate this equation on labor supply:
@@
$Y_{ipct}$ $=\alpha_{g}\;+\;\beta_{1g}\bigl[\mathrm{SSP}_{ipct}\times\mathrm{ENDA}_{ct}\bigr]\;+\;\lambda_{s(c),g}\bigl[\mathrm{SSP}_{ipct}\times\mathrm{SSM}_{s(c),t}\bigr]$ (1)
$\quad+\theta_{gct}+\mu_{gpc}+\kappa_{gpt}+X_{ipct}^{\prime}\,\delta_{g}+\epsilon_{ipct}.$
@@
In our regression, equation (1) defines the labor market outcomes $Y_{ipct}$ for individual $i$ of sex $g$, in partnership type $p$, residing in county $c$ in state $s$, at time $t$. These outcomes include labor force participation, employment status, and the percentile rank of real wages (in 2019 dollars). The binary variable $\mathrm{SSP}_{ipct}$ indicates whether an individual is in a same‐sex partnership (1) or a different‐sex partnership (0). $\mathrm{SSM}_{s(c),t}$ indicates whether the state $s(c)$ recognizes same‐sex marriage in year $t$, where $s(c)$ denotes the state corresponding to county $c$. Our key coefficient of interest is $\beta_{1g}$, which measures the sex-specific effect of the interaction of same‐sex partnership status with the presence of a county‐level Employment Non‐Discrimination Act, $\mathrm{ENDA}_{ct}$. Treatment status is computed using the earliest year of ENDA implementation at the county or state level, with individuals remaining classified as “treated” thereafter. All coefficients ($\alpha_{g},\beta_{1g},\lambda_{s(c),g},\delta_{g}$) and fixed effects ($\theta_{gct},\mu_{gpc},\kappa_{gpt}$) are fully interacted with sex $g$, so that the model is estimated jointly but allows each sex to have its own coefficients. The error term $\epsilon_{ipct}$ is clustered at the county level to account for within-county correlation. [^8] The ACS only identifies metro counties, so the county fixed effect is close to a pseudo-county fixed effect. Rural counties within a state are compared together with a true county fixed effect for metro area counties.
We restrict the sample to individuals aged 25–65 and estimate all models jointly by sex using ACS person weights. Following chen2024logs, we compute weighted within-year percentile ranks for real annual earnings and real hourly wages (expressed in 2019 dollars) by assigning individuals to weighted percentiles each year. We then construct three distributional indicators, at the 25th, 50th, and 75th percentiles, set to one when an individual’s percentile rank meets or exceeds each threshold. Because ACS earnings reflect a 23-month reference period (hansen2020labor), measured earnings may not align exactly with ENDA adoption dates, likely attenuating treatment effects.
@@
We include three sets of high-dimensional fixed effects, each fully interacted with sex to reflect the joint estimation. First, the county-by-year fixed effects $\theta_{gct}$ absorb all time-varying shocks or policy changes common to individuals of sex $g$ within county $c$ and year $t$. Second, the partnership-by-county fixed effects $\mu_{gpc}$ capture time-invariant characteristics specific to partnership type $p$ within county $c$ for each sex $g$. Third, the partnership-by-year fixed effects $\kappa_{gpt}$ absorb year-specific shocks that differentially affect partnership types for each sex. The vector $X_{ipct}^{\prime}\delta_{g}$ includes demographic controls (race, age, and education), also allowed to vary by sex.
The coefficient of interest, $\beta_{1g}$, identifies how ENDA coverage $\mathrm{ENDA}_{ct}$ differentially affects same-sex partners relative to different-sex partners for each sex $g$, net of any confounding from same-sex marriage legalization. The interacted term $\lambda_{s(c),g}\bigl[\mathrm{SSP}_{ipct}\times\mathrm{SSM}_{s(c),t}\bigr]$ partials out state-specific labor-market impacts of same-sex marriage legalization on same-sex partners for each sex.
The identifying variation for $\beta_{1g}$ comes from differences in the timing of county-level ENDA adoption, combined with how these policy changes differentially affect individuals in same-sex versus different-sex couples within counties, holding constant sex-specific fixed effects. This allows us to isolate how employment protections change labor supply and wage outcomes for same-sex partners relative to different-sex partners of the same sex. Thus, $\beta_{1g}$ captures the sex-specific differential impact of ENDA on labor market outcomes for individuals in same-sex couples. [^9] We recognize that the effects of anti-discrimination laws may vary across regions (e.g., Burn2018), but our objective is to estimate an average effect for those directly subject to the laws. To interpret $\beta_{1g}$ as the causal effect of enacting employment protections for sexual orientation, we assume that no other factors differentially shift outcomes for same-sex versus different-sex partners in counties that do and do not adopt ENDAs. Below, we outline the threats to this assumption and describe how our empirical design addresses each one.
3.1 Threats to Identification
We group the main identification concerns into three categories: (i) threats that can be addressed directly through sample restrictions, (ii) threats mitigated through design choices and high-dimensional fixed effects, and (iii) residual concerns that cannot be fully ruled out but primarily generate attenuation rather than spurious effects.
3.1.1 Threats Addressed
Selective Migration.
Individuals in same-sex couples may relocate to jurisdictions with stronger LGBT protections, altering the composition of treated counties. To assess this possibility, we augment the specification with an additional layer of interactions distinguishing non-mover females, non-mover males, mover females, and mover males. This four-way structure lets ENDA effects differ flexibly by migration status. Comparing coefficients across these groups provides a direct test for migration-induced bias, and the results show no meaningful differences. [^10] Long-run unobserved mobility may still shift the composition of same-sex couples in treated areas, though our mover–non-mover comparisons indicate limited bias.
Household Formation and Sample Composition.
A potential concern is that ENDAs may affect the likelihood that individuals enter or remain in a same-sex partnership, thereby altering who appears in our coupled-household sample. If anti-discrimination protections increase relationship stability or the willingness to publicly identify a partner, then treatment could shift the composition of the same-sex population in ways correlated with labor market outcomes. To assess this concern, we estimate whether the share of adults living in same-sex couples changes following ENDA adoption at the county level. We find no evidence of differential changes for men or women. This suggests that ENDAs do not meaningfully affect partnership formation or household structure, and thus do not induce endogenous entry into the coupled sample. As a result, the identifying variation in our main specification is unlikely to reflect compositional changes in who is observed as part of a same-sex household.
Residence–Work Geography Mismatch.
ENDAs apply to the county of employment, whereas ACS geography reflects the county of residence. As with migration, this creates the possibility of misclassifying some individuals with respect to treatment if they commute across county (or state) lines. We evaluate this concern using the same strategy as our mover–non-mover analysis: we allow ENDA effects to differ flexibly across groups likely to experience geographic mismatch. The estimates are stable across these specifications, indicating that any residence–work misalignment does not materially influence the results. As with other sources of geographic measurement error, this type of misclassification primarily generates attenuation rather than spurious treatment effects.
3.1.2 Threats Mitigated
Staggered Adoption.
To address potential biases from staggered policy adoption, we employ the methodology of borusyak2021revisiting. Unlike standard TWFE models, which may misattribute post-treatment effects from early-treated units to later-treated ones, their estimator uses only untreated observations to form counterfactuals for treated units. Because we estimate the model jointly by sex, this approach is applied separately within each sex group, yielding the sex-specific treatment effects $\beta_{1g}$ without contamination from negative weighting. Our identification hinges on the parallel trends assumption: in the absence of ENDAs, labor market trends for same-sex and different-sex partners would have evolved similarly within each sex group across treated and untreated counties. [^11] This requires no anticipatory responses and equal pre-treatment trends within sex-by-partnership cells across treatment timing groups. We assess this assumption with event-study analyses (Figures 2–3), which show no evidence of differential pre-trends. [^12] We conduct a sensitivity analysis following RambachanRoth2023. As detailed in Appendix Table C.8 and discussed in Appendix C.7, we find that our extensive margin estimates for men are robust to post-treatment trend violations up to 40% of the maximal pre-trend deviation, while the null results for women persist even under relaxed assumptions.
Same-Sex Marriage Legalization.
Same-sex marriage (SSM) legalization may confound the effects of ENDAs by shifting household specialization or labor supply among same-sex partners. We directly partial out these effects by including sex-specific interactions of same-sex partnership status with SSM legalization status. This absorbs any differential impacts of state-level marriage reforms on same-sex couples. [^13] While these controls mitigate potential bias, overlapping policies may still exert residual confounding influences on the estimated effects. As an additional check, Section 5.2 restricts the sample to pre-SSM years within each state, again yielding similar estimates.
Differential Local Trends and Endogenous Adoption.
Jurisdictions adopting ENDAs tend to be more supportive toward sexual minorities. If positive sentiment drives both ENDA passage and improvements in labor market outcomes, estimates may reflect omitted variables. Our saturated triple-difference specification addresses this by including county-year, partnership-by-year, and partnership-by-county fixed effects, all interacted with sex. These fixed effects ensure comparisons are made only within county-year cells and within sex-by-partnership cells, substantially reducing concerns about correlated local shocks or endogenous policy timing.
3.1.3 Residual Threats
Misclassification of Sexual Orientation.
Our comparison relies on household composition rather than self-identified sexual orientation. Bisexual individuals in different-sex relationships enter the control group, and some gay or lesbian workers may also be misclassified into that group. Because lesbian women tend to earn more than heterosexual women, gay men tend to earn less than heterosexual men, and bisexual workers earn less than both (mize2016sexual), such misclassification does not shift group earnings in a uniform direction. In addition, bisexual workers may respond differently to ENDAs than gay or lesbian workers, so the sign of any resulting bias in the estimated ENDA coefficients is theoretically ambiguous. At the same time, misclassification that places gay or lesbian individuals into the control group reduces the contrast between partnership types and therefore operates through the usual attenuation mechanism. Overall, these forms of misclassification complicate interpretation but are unlikely to generate spurious positive treatment effects.
Overlap and Heterogeneity in Local and State ENDA Laws.
Local and state ENDAs vary in scope and enforcement. Our coding focuses on the first year in which a county receives broad protections in private employment and housing. For example, although Delaware extended protections to state employees in 1999, we code treatment beginning in 2009 when comprehensive protections were enacted. Similarly, some local ordinances initially covered only public employees before broader laws passed. These earlier, narrower protections may still have influenced social attitudes or employer behavior prior to full coverage. In addition, the strength of protections differs across jurisdictions, some laws exempt small firms or specific sectors, and many large employers adopt nondiscrimination policies independent of legal requirements (e.g., firms scoring highly on the Corporate Equality Index). Collapsing this heterogeneity into a single binary ENDA indicator necessarily compresses meaningful variation. This generates classical attenuation bias, pushing estimated effects toward zero and making our findings conservative.
Geographic Misclassification From PUMA-to-County Assignment.
ACS microdata identify geography at the PUMA level, which sometimes maps imperfectly to counties. A small share of individuals may thus be assigned to counties with classification error. We do not exclude ambiguous PUMAs; however, such misclassification behaves like classical measurement error in the treatment indicator and therefore attenuates estimated treatment effects rather than creating false positives.
To summarize, our empirical strategy exploits a fully saturated triple-differences design estimated jointly by sex. We control for confounding policy changes such as same-sex marriage legalization using sex-specific marriage interactions, and we implement the estimator of borusyak2021revisiting to mitigate biases arising from staggered ENDA adoption. Together, these features allow us to identify how local ENDAs affect the labor market outcomes of individuals in same-sex couples relative to their different-sex counterparts within each sex group. As we show in Appendix C.1, the estimates for males and females are similar in sign and magnitude across alternative estimators, supporting the robustness of our findings.